A gaseous compound has a density of 1.21 g/L at 23.4 degree C and 1.24 atm. What is the molar mass of the compound?
 
Molar mass = _______ 
Answer #1

PV = nRT

n = m/MW

PV = m/MW*RT

m/V =D = P*MW/(RT)

MW = D*RT/P

MW = (1.21)(0.082)(23.4+273)/1.24 = 23.7167g /mol